OpenStudy (anonymous):

9=90,8=72,7=56,6=42,3=?

OpenStudy (anonymous):

As 9=90=9*10 8=72=8*9 7=56=7*8 6=42=6*7 so we can find that x=x*(x+1). Thus 3=3*4=12

OpenStudy (anonymous):

dalbert?